OpenStudy (anonymous):

-1 3/7 x (-3 2/3)?? Can someone please help me?

OpenStudy (anonymous):

Hello, Can someone please help me?

OpenStudy (goldphenoix):

\[\large \large -1 \frac{ 3 }{ 7 } \times -3\frac{ 2 }{ 3 }\] Does your question look like that?

OpenStudy (anonymous):

yes but the ( -3 2/3 )